# KIAS-Q25002

### v2

@@ 1 Introduction
in terms of the amplitude $T(\pi)$ of the Goldstone boson $\pi$, where the signs indicate the incoming (--) or outgoing ($+$) gauge boson $A$, with momentum $p$. This formula suggests that currents associated with massive gauge bosons exhibit modified conservation properties in the four-dimensional framework. Given that the currents associated with massless gauge bosons play a crucial role in ensuring the full absence of non-physical momentum dependence, it would be beneficial to preserve current conservation even for massive gauge bosons in order to identify non-physical contributions by adopting the 5V framework.
@@
After the introduction, Section 2 derives the 5V $R_{\xi}$ gauge-fixing term in the Lagrangian of the Abelian Higgs model with the complex Higgs field in the polar basis, obtained by imposing BRST invariance on the Lagrangian including the gauge-fixing term. Subsequently, the three physical polarization vectors, which are transverse to the five-momentum, are determined by analyzing the state operators and identifying the non-physical ones that do not belong to the cohomology. Afterward, the physical implications of the third transverse mode, corresponding to the longitudinal mode in the 4V description, are examined in detail. Section 3 examines the tree-level amplitudes for the processes $A_{L}A_{L}\rightarrow A_{L}A_{L}$ and $A_{L}A_{L}\rightarrow hh$, involving massive gauge bosons $A_{L}$ with longitudinal polarization and Higgs bosons $h$. It is shown that the tree-level amplitudes for each Feynman diagram, derived from the 5V $R_{\xi}$ gauge, are equivalent to those from the 4V $R_{\xi}$ and FD gauges in the polar basis. Section 4 presents the high-energy divergences in the amplitudes through explicit computations and identifies their origin from the equivalence between the high-energy and vanishing VEV limits, ensuring no gauge cancellation in the Cartesian basis for the complex Higgs field. After discussing the rationale for the high-energy divergences in the 4V $R_{\xi}$ gauge and the origin of no gauge cancellation in the FD gauge, the impact of the 5V $R_{\xi}$ gauge is explored through a rigorous comparison of the amplitudes from the 5V $R_{\xi}$, 4V $R_{\xi}$, and FD gauges. Section 5 summarizes the key findings and examines the physical implications of the 5V framework. The ingredients for the amplitude calculations are provided in Appendix A.
@@
The Lagrangian in Eq. (4) expressed in the 5V description is invariant under the gauge transformation $A^{M}\rightarrow A^{M}+\partial^{M}\Lambda$.
@@
$\mathcal{G}(A)=\partial_{M}A^{M},$ (15)
resulting in the combination to be given by
$\mathcal{L}_{\scriptsize\mbox{gf+gh}}=-\frac{(\partial_{M}A^{M})^{2}}{2\xi}-\bar{c}(\partial^{M}\partial_{M})c.$ (16)
@@
$\mathcal{L}_{\scriptsize\mbox{free}}=-\frac{1}{4}F^{MN}F_{MN}-\frac{(\partial_{M}A^{M})^{2}}{2\xi}.$ (17)
Employing this expression, we can find the Fourier-transformed EOM in Eq. (17) in the momentum space as
$O^{M}_{\;\;\,N}\tilde{A}^{N}(P)=0\quad\mbox{with}\quad O^{M}_{\;\;\,N}=-(P^{R}P_{R}^{*})\,g^{M}_{\;\;\,N}+P^{M}P_{N}^{*}-\frac{P^{M}P^{*}_{N}}{\xi},$ (18)
@@
$c$ $=\int\frac{d^{3}\vec{p}}{(2\pi)^{3}2E}\Big{[}C(p)e^{-ip\cdot x}+C^{\dagger}(p)e^{ip\cdot x}\Big{]},$ (24)
$\bar{c}$ $=\int\frac{d^{3}\vec{p}}{(2\pi)^{3}2E}\Big{[}\bar{C}(p)e^{-ip\cdot x}+\bar{C}^{\dagger}(p)e^{ip\cdot x}\Big{]},$ (25)
@@
in terms of the polar $\theta$ and azimuthal $\phi$ angles. Here, each polarization vector is orthogonal to other ones for the 5V inner product. Especially, the space-like polarization vectors, ϵ(±)italic-$\epsilon(\pm)$ and $\epsilon(0)$, satisfies the following orthonormality,
@@
[…]
which are valid even in the massless limit ($m\to 0$). These relations allow us to eliminate the unphysical polarization modes corresponding to the light-like vectors with indices $>$ and $<$ in the context of cohomology; any physical state must be in the cohomology [9, 10]. Let us consider a normalized physical state $|\Psi\rangle$ satisfying $\langle\Psi|\Psi\rangle=1$ and $Q_{B}|\Psi\rangle=0$ in the cohomology. The normalization condition implies that $|\Psi\rangle\neq Q_{B}|\Psi^{\prime}\rangle$ for another physical state $|\Psi^{\prime}\rangle$, due to the nilpotent nature of the BRST operation $Q_{B}^{2}=0$, with the Hermitian operator $Q_{B}$, resulting in non-normalizability of the state $|\Psi\rangle$. Eq. (34) ensures that the state $a^{\dagger}_{>}(p)|\Psi\rangle$ is not in the cohomology because the BRST operation $Q_{B}a^{\dagger}_{>}(p)|\Psi\rangle=\sqrt{2}EC^{\dagger}(p)|\Psi\rangle$ on this state does not yield zero. Eq. (35) allows us to express the state $a^{\dagger}_{<}(p)|\Psi\rangle$ as a combination of two states: $Q_{B}\bar{C}(p)|\Psi\rangle$ and $a^{\dagger}_{>}(p)|\Psi\rangle$. This combination also does not vanish under the BRST operation, implying that the state is not in the cohomology. In the massless limit, the state $a^{\dagger}_{<}(p)|\Psi\rangle$ can become proportional to the state $Q_{B}\bar{C}(p)|\Psi\rangle$, which prevents the state from being in the cohomology due to its non-normalizability.
@@ 4 Identification of the high-energy divergence in tree-level amplitudes
respectively, for a gauge boson moving along the $z$ axis, with the on-shell energy $E=\sqrt{p^{2}+m^{2}}$. For the 4V $R_{\xi}$ gauge, the polarization vector $\epsilon^{M}_{0}(0)$, with an additional five-momentum factor added to $\epsilon^{M}(0)$, causes high-energy divergences in each amplitude because the interaction vertices in the Cartesian basis are not divergence-free, resulting in the non-physical five-momentum contribution to each amplitude. For the FD gauge, although there is an additional five-momentum factor, it vanishes in the high-energy limit, resulting in no gauge cancellation, similar to the 5V $R_{\xi}$ gauge.
@@
For the process $A_{L}A_{L}\rightarrow A_{L}A_{L}$, only the 4V $R_{\xi}$ gauge exhibits amplitudes significantly larger than their sum. The amplitudes from the FD gauge show slightly different behavior compared to those of the 5V $R_{\xi}$ gauge, which originates from the additional 5V momentum factor in Eq. (73). Similarly, for the $A_{L}A_{L}\rightarrow hh$ process, only the 4V $R_{\xi}$ gauge exhibits amplitudes significantly larger than their sum, and the FD gauge amplitudes deviate slightly from those of the 5V $R_{\xi}$ gauge, but neither the 5V $R_{\xi}$ gauge nor the FD gauge results in gauge cancellation.
@@ 5 Conclusion
Next, we showed that the 5V framework can straightforwardly trace the origin of the high-energy divergences in the amplitudes from the 4V $R_{\xi}$ gauge, which arises from the non-physical five-momentum contribution. While the FD gauge also includes a five-momentum contribution, it vanishes in the high-energy limit, resulting in no gauge cancellation, similar to the 5V $R_{\xi}$ gauge. These findings were explicitly demonstrated through a quantitative investigation of the amplitudes and their sums.
The structure of the third transverse polarization vector of the gauge boson is expected to deepen our understanding of the dynamics of massive gauge bosons. The simple analytic structure of the 5V framework, along with the absence of gauge cancellation in the 5V $R_{\xi}$ gauge, provides a useful criterion for assessing the degree of high-energy divergences in tree-level amplitudes across various gauges. Thus, this approach offers significant practical advantages for both analytical and numerical amplitude calculations. Notably, this work represents the first extension of defining the asymptotic on-shell gauge boson state with the Goldstone state, where the kinematic factor is not a constant [2, 3]. This observation suggests that a natural next step is to generalize the gauge boson state with an arbitrary kinematic factor at the Goldstone state. Furthermore, this framework can be extended to non-Abelian gauge theories, which incorporate the self-interactions of gauge bosons, thereby paving the way for even broader applications. This extension is currently under investigation and will be reported separately, with potential implications for a wider class of gauge theories.
@@ Acknowledgments
We thank Prof. Andrea Wulzer, Prof. Alex Pomarol, and Dr. Kiyoharu Kawana for their valuable comments on this work. This work was supported by a KIAS Individual Grant (QP090001) through the Quantum Universe Center at the Korea Institute for Advanced Study. We also acknowledge the Institut de Física d’Altes Energies (IFAE) for their hospitality during our visit.

### v3

@@ 1 Introduction
in terms of the amplitude $T(\pi)$ of the Goldstone boson $\pi$, where the signs indicate the incoming (--) or outgoing ($+$) gauge boson $A$, with momentum $p$. This formula suggests that currents associated with massive gauge bosons exhibit modified conservation properties in the four-dimensional framework. Given that the currents associated with massless gauge bosons play a crucial role in ensuring the full absence of unphysical momentum dependence, it would be beneficial to preserve current conservation even for massive gauge bosons in order to identify unphysical contributions by adopting the 5V framework.
@@
After the introduction, Section 2 derives the 5V $R_{\xi}$ gauge-fixing term in the Lagrangian of the Abelian Higgs model with the complex Higgs field in the polar basis, obtained by imposing BRST invariance on the Lagrangian including the gauge-fixing term. Subsequently, the three physical polarization vectors, which are transverse to the five-momentum, are determined by analyzing the state operators and identifying the unphysical ones that do not belong to the cohomology. Afterward, the physical implications of the third transverse mode, corresponding to the longitudinal mode in the 4V description, are examined in detail. Section 3 examines the tree-level amplitudes for the processes $A_{L}A_{L}\rightarrow A_{L}A_{L}$ and $A_{L}A_{L}\rightarrow hh$, involving massive gauge bosons $A_{L}$ with longitudinal polarization and Higgs bosons $h$. It is shown that the tree-level amplitudes for each Feynman diagram, derived from the 5V $R_{\xi}$ gauge, are equivalent to those from the 4V $R_{\xi}$ and FD gauges in the polar basis. Section 4 presents the high-energy divergences in the amplitudes through explicit computations and identifies their origin from the equivalence between the high-energy and vanishing VEV limits, ensuring no gauge cancellation in the Cartesian basis for the complex Higgs field. After discussing the rationale for the high-energy divergences in the 4V $R_{\xi}$ gauge and the origin of no gauge cancellation in the FD gauge, the impact of the 5V $R_{\xi}$ gauge is explored through a rigorous comparison of the amplitudes from the 5V $R_{\xi}$, 4V $R_{\xi}$, and FD gauges. Section 5 summarizes the key findings and examines the physical implications of the 5V framework. The ingredients for the amplitude calculations are provided in Appendix A.
@@
Note that the Lagrangian in Eq. (4) expressed in the 5V description is invariant under the gauge transformation $A^{M}\rightarrow A^{M}+\partial^{M}\Lambda$.
@@
$\mathcal{G}(A)=\bar{\partial}_{M}A^{M},$ (15)
where ¯$\bar{\partial}^{M}=(\partial^{\mu},-m)=\partial_{M}$, referred to as the conjugated 5V derivative. This results in the combination to be given by
$\mathcal{L}_{\scriptsize\mbox{gf+gh}}=-\frac{(\bar{\partial}_{M}A^{M})^{2}}{2\xi}-\bar{c}(\bar{\partial}^{M}\partial_{M})c.$ (16)
@@
$\mathcal{L}_{\scriptsize\mbox{free}}=-\frac{1}{4}F^{MN}F_{MN}-\frac{(\bar{\partial}_{M}A^{M})^{2}}{2\xi}.$ (17)
Employing this expression, we can find the free-theory EOM as
$\begin{pmatrix}(\square+m^{2})g^{\mu}_{\;\,\nu}-\Big{(}1-\frac{1}{\xi}\Big{)}\partial^{\mu}\partial_{\nu}&-m\Big{(}1-\frac{1}{\xi}\Big{)}\partial^{\mu}\\ -m\Big{(}1-\frac{1}{\xi}\Big{)}\partial_{\nu}&\square+\frac{m^{2}}{\xi}\end{pmatrix}\begin{pmatrix}A^{\mu}(x)\\ A^{4}(x)\end{pmatrix}=0.$ (18)
Fourier transforming to momentum space leads to
$O^{M}_{\;\;\,N}\tilde{A}^{N}(P)=0\quad\mbox{with}\quad O^{M}_{\;\;\,N}=-(P^{R}P_{R}^{*})\,g^{M}_{\;\;\,N}+\bigg{(}1-\frac{1}{\xi}\bigg{)}P^{M}P_{N}^{*},$ (19)
@@
$c$ $=\int\frac{d^{3}\vec{p}}{(2\pi)^{3}\sqrt{2E}}\Big{[}C(p)e^{-ip\cdot x}+C^{\dagger}(p)e^{ip\cdot x}\Big{]},$ (25)
$\bar{c}$ $=\int\frac{d^{3}\vec{p}}{(2\pi)^{3}\sqrt{2E}}\Big{[}\bar{C}(p)e^{-ip\cdot x}+\bar{C}^{\dagger}(p)e^{ip\cdot x}\Big{]},$ (26)
@@
in terms of the polar $\theta$ and azimuthal $\phi$ angles. Here, the space-like polarization vectors, ϵ(±)italic-$\epsilon(\pm)$ and $\epsilon(0)$, satisfy the following orthonormality,
@@
$\{Q_{B},\bar{C}^{\dagger}(p)\}$ $=-\xi^{-1}\sqrt{2}E\,a^{\dagger}_{<}(p),$ (36)
which are valid even in the massless limit ($m\to 0$). These relations allow us to eliminate the unphysical polarization modes corresponding to the light-like vectors with indices $>$ and $<$ in the context of cohomology; any physical state must be in the cohomology [9, 10]. A physical state $|\Psi\rangle$ defined as a normalized and BRST-closed state, satisfying $\langle\Psi|\Psi\rangle=1$ and $Q_{B}|\Psi\rangle=0$. The normalization condition implies that a physical state $|\Psi\rangle$ cannot be written in the BRST-exact form $Q_{B}|\Psi^{\prime}\rangle$ for any state $|\Psi^{\prime}\rangle$, since such a state has vanishing norm as a consequence of the nilpotency $Q_{B}^{2}=0$ of the Hermitian BRST operator $Q_{B}$. According to Eq. (35), the state $a^{\dagger}_{>}(p)|\Psi\rangle$ is not BRST-closed, since it is mapped by $Q_{B}$ to a nonzero state: $Q_{B}a^{\dagger}_{>}(p)|\Psi\rangle=\sqrt{2}E\,C^{\dagger}(p)|\Psi\rangle$. On the other hand, Eq. (36) shows that $a^{\dagger}_{<}(p)|\Psi\rangle$ is proportional to $Q_{B}\bar{C}(p)|\Psi\rangle$, implying that it is BRST-exact and thus does not belong to the cohomology.
@@ 4 Identification of the high-energy divergence in tree-level amplitudes
respectively, for a gauge boson moving along the $z$ axis, with the on-shell energy $E=\sqrt{p^{2}+m^{2}}$. For the 4V $R_{\xi}$ gauge, the polarization vector $\epsilon^{M}_{0}(0)$, with an additional five-momentum factor added to $\epsilon^{M}(0)$, causes high-energy divergences in each amplitude because the interaction vertices in the Cartesian basis are not divergence-free, resulting in the unphysical five-momentum contribution to each amplitude. For the FD gauge, although there is an additional five-momentum factor, it vanishes in the high-energy limit, resulting in no gauge cancellation, similar to the 5V $R_{\xi}$ gauge.
@@
For the process $A_{L}A_{L}\rightarrow A_{L}A_{L}$, only the 4V $R_{\xi}$ gauge exhibits amplitudes significantly larger than their sum. The amplitudes from the FD gauge show slightly different behavior compared to those of the 5V $R_{\xi}$ gauge, which originates from the additional five-momentum factor in Eq. (74). Similarly, for the $A_{L}A_{L}\rightarrow hh$ process, only the 4V $R_{\xi}$ gauge exhibits amplitudes significantly larger than their sum, and the FD gauge amplitudes deviate slightly from those of the 5V $R_{\xi}$ gauge, but neither the 5V $R_{\xi}$ gauge nor the FD gauge results in gauge cancellation.
@@ 5 Conclusion
Next, we showed that the 5V framework can straightforwardly trace the origin of the high-energy divergences in the amplitudes from the 4V $R_{\xi}$ gauge, which arises from the unphysical five-momentum contribution. While the FD gauge also includes a five-momentum contribution, it vanishes in the high-energy limit, resulting in no gauge cancellation, similar to the 5V $R_{\xi}$ gauge. These findings were explicitly demonstrated through a quantitative investigation of the amplitudes and their sums.
The structure of the third transverse polarization vector of the gauge boson is expected to deepen our understanding of the dynamics of massive gauge bosons. The simple analytic structure of the 5V framework, along with the absence of gauge cancellation in the 5V $R_{\xi}$ gauge, provides a useful criterion for assessing the degree of high-energy divergences in tree-level amplitudes computed in various gauges. Thus, this approach offers significant practical advantages for both analytical and numerical amplitude calculations. Notably, this work presents the first extension in which the asymptotic on-shell gauge boson state is defined together with a Goldstone state multiplied by a non-constant kinematic factor [2, 3]. This observation suggests that a natural next step is to generalize the gauge boson state with an arbitrary kinematic factor at the Goldstone state. Furthermore, this framework can be extended to non-Abelian gauge theories, which incorporate the self-interactions of gauge bosons, thereby paving the way for even broader applications. This extension is currently under investigation and will be reported separately, with potential implications for a wider class of gauge theories.
@@ Acknowledgments
We thank Prof. Andrea Wulzer, Prof. Alex Pomarol, Dr. Kiyoharu Kawana, and Dr. Yu-hui Zheng for their valuable comments on this work. This work was supported by a KIAS Individual Grant (QP090001) through the Quantum Universe Center at the Korea Institute for Advanced Study. We also acknowledge the Institut de Física d’Altes Energies (IFAE) for their hospitality during our visit.